# multiple directions/.style=for tree=#1, phantom, for relative level=1no edge, delay=!c.content/.pgfmath=content("!u"), before computing xy=l=0,s=0, multiple directions/.default=, grow subtree/.style=for tree=grow=#1, grow’ subtree/.style=for tree=grow’=#1

### v1

@@ 3.4 Uniqueness
Lemma 10 implies that $\mathbb{P}_{Z\sim\zeta_{h}}(Z=0)>0$. This uniform lower bound does not depend on the values of $w(v_{0},v_{-1}),Z_{h}(v_{0},v_{-1})$, so it is $\mathcal{P}$ almost-sure. ∎
@@
$u_{1}(u)=\operatorname*{argmax}_{y\sim u}(w^{s}(u,y)-Z_{h}^{s}(u,y)),u_{2}(u)=\operatorname*{\overset{[2]}{argmax}}_{y\sim u_{1}}(w^{s}(u_{1},y)-Z_{h}^{s}(u_{1},y)),$
@@
$u_{2n+1}(u)=\operatorname*{argmax}_{y\sim u_{2n}}(w^{s}(u_{2n},y)-Z_{h}^{s}(a_{2n},y)),u_{2n+2}(u)=\operatorname*{\overset{[2]}{argmax}}_{y\sim u_{2n+1}}(w^{s}(u_{2n+1},y)-Z_{h}^{s}(u_{2n+1},y)),$
@@
4 From infinite to finite
In this entire section, unless stated otherwise, we will adopt the edge-rooted point of view.
The aim of this section is to prove the following statement:
Proposition 16.
Assume $(G_{n},o_{n})_{n\geq 1}$ a sequence of finite graphs which converges locally to the unimodular BGW tree $(\mathbb{T},o)$ with reproduction law $\pi$ and weights $\omega$.
@@ 4 From infinite to finite
The proof can be broken down into two parts, one easy, and one hard:
@@
For the easy part, using tightness arguments, we will show that every subsequence has a convergent subsequence. The limit of any convergent subsequence is necessarily suboptimal by the definition of optimality, one only needs to verify that the limit will define a unimodular matched graph.
For the hard part, we will show that the limit of any convergent subsequence is necessarily optimal by constructing an almost optimal quasi-matching on $G_{n}$ from the optimal matching $(\mathbb{T},\mathbb{M}_{h})$. Indeed, as a consequence, this will show that $M_{\mathrm{opt}}(G_{n})$ does asymptotically better than $\operatorname{perf_{V}}(\mathbb{M}_{h})$.
@@ 4.1 Local convergence of subsequences
Lemma 17.
Let $(G_{n},o_{n})$ be a sequence in the space of rooted graphs, $G_{n}=(V_{n},E_{n})$. Let $(f_{n})_{n\in\mathbb{N}}$ be any (random) functions on $E_{n}$ such that the sequence $(f_{n})$ is uniformly bounded. Assume that $G_{n}$ converges locally to an almost surely locally finite graph $G$. Then for any $H\in\mathbb{N}$, the sequence of decorated graphs $N_{H}(G_{n},o_{n},f_{n})$ is tight.
@@
Let $H\in\mathbb{N}$ and $\varepsilon>0$. Denote by $\tau_{H,N}$ the set of rooted graphs of depth $H$ and degree bounded by $N$. Since $G$ is locally finite, for $N$ large enough, we have $\mathbb{P}(N_{H}(G,o)\in\tau_{H,N})>1-\frac{\epsilon}{2}$. By local convergence, $\mathbb{P}(N_{H}(G_{n},o_{n})\in\tau_{H,N})\underset{n\rightarrow\infty}{\longrightarrow}\mathbb{P}(N_{H}(G,o)\in\tau_{H,N})$ hence for $n$ large enough,
$\mathbb{P}(N_{H}(G_{n},o_{n})\in\tau_{H,N})>1-\epsilon$
@@
$\mathbb{P}(N_{H}(G_{n},o_{n},f_{n})\in\tau_{H,N,A})>1-\epsilon,$
$N_{H}(G_{n},o,f_{n})$ is thus tight. ∎
We can now go back to the proof of Proposition 16. Take $f_{n}=\mathbbm{1}_{M_{\mathrm{opt}}(G_{n})}$, it is uniformly bounded by $1$. Apply Lemma 17 to $N_{0}(G_{n},o_{n},f_{n})$ to obtain a subsequence such that $N_{0}(G_{k_{n}},o_{k_{n}},f_{k_{n}})$ converges.
But $(G_{k_{n}},o_{k_{n}})$ still converges locally to $(G,o)$, so apply Lemma 17 again to find a subsequence $k_{n}^{\prime}$ of $k_{n}$ such that $N_{1}(G_{k^{\prime}_{n}},o_{k^{\prime}_{n}},f_{k^{\prime}_{n}})$ converges locally. Repeat the process and use a diagonal argument to obtain a subsequence $l_{n}$ such that for all $H\in\mathbb{N}$, $N_{H}(G_{l_{n}},o_{l_{n}},f_{l_{n}})$ converges weakly to some limit $N_{H}(G,o,g_{H})$ for some decoration $g_{H}$.
Furthermore, since the decorations $g_{H}$ are obtained from a diagonal construction, the family $(g_{H})_{H\geq 0}$ is consistent and therefore define a limiting decoration $g$. To prove that this decoration is a matching, we just have to take successively limits in $n\rightarrow\infty$ and $H\rightarrow\infty$ in the equalities
$\mathbb{P}\left(\sum_{u\sim v}f_{l_{n}}(u,v)\leq 1,(u,v)\in V(N_{H}(G_{n},o_{n}))\right)=1,\mathbb{P}\left(f_{l_{n}}(e)\in\{0,1\},e\in E(N_{H}(G_{n},o_{n}))\right)=1.$
@@ 4.2 Optimality of limits of subsequences
In this section, we carry out the hard part of the proof of Proposition 16. We want to reconstruct a quasi matching on $G_{n}=(V_{n},E_{n})$ from $(\mathbb{T},\mathbb{M}_{h})$. The crucial idea is, for every $(i,j)\in V_{n}^{2}$, to compute the probability that $(i,j)\in\mathbb{M}_{h}$ conditionally on the $H$-neighbourhood of $o$ in $\mathbb{T}$ being the corresponding one of $(i,j)$ in $G_{n}$. Informally, this constructs a "score" matrix where each score is the likelihood of matching $(i,j)$ by looking up to depth $H$.
The first hurdle is that the $H$- neighbourhood of $(i,j)$ in $G_{n}$ needs not to be a tree, so we will consider the universal cover of $G_{n}$ rooted at $(i,j)$ instead, and control the resulting error.
@@
By construction, this random matching on $G_{n}$ will have the same performance as $(\mathbb{T},\mathbb{M}_{h})$ with asymptotically small error, which concludes.
@@ 4.2.1 Score function on edges of finite graphs
Let $(i,j)\in V_{n}^{2}$, we define $G_{n,\infty}^{(i,j)}$ as $\emptyset$ if $(i,j)$ is not an edge of $G_{n}$, and as the universal cover $G_{n,\infty}^{(i,j)}$ of $G_{n}$ rooted at $(i,j)$ if $(i,j)$ is an edge of $G_{n}$. This cover is the tree of non-backtracking trajectories from the edge $(i,j)$. Namely, when $(i,j)$ is an edge of $G_{n}$, $G_{n,\infty}^{(i,j)}$ is a tree rooted at the edge $(i,j)$, children of $j$ children of $i$ (resp. $j$) are the neighbours of $i$ (resp. $j$) in $G_{n}$, with $j$ (resp. $i$ excluded). The children of any vertex $v$ with parent $v^{\prime}$ are then the neighbours in $G_{n}$ of $v$, with $v^{\prime}$ excluded. See Figure 14 for an illustration. For $H\geq 0$, we set $G_{n,H}^{(i,j)}:=N_{H}(G_{n,\infty}^{(i,j)})$.
@@
We will need the following that translates the fact that the universal cover of $G_{n}$ converges locally to the same limit as $G_{n}$.
Lemma 18.
Let $(i,j)$ be chosen uniformly in $(V_{n})^{2}$, then:
$G_{n,\infty}^{(i,j)}|(i,j)\in E_{n}\underset{n\rightarrow\infty}{\overset{loc}{\longrightarrow}}\mathbb{T}.$
Proof of Lemma 18.
@@
$\mathbb{P}\left(N_{H}(G_{n},(i,j))|(i,j)\in E_{n},(i,j)\text{ uniformly chosen})\text{ is a tree}\right)\underset{n\rightarrow\infty}{\longrightarrow}\mathbb{P}\left(N_{H}(\mathbb{T})\text{ is a tree}\right)=1.$
So w.h.p $N_{H}((G_{n},(i,j)))$ is a tree, in which case the universal cover of $G_{n}$ rooted on $o_{n}$ up to height $H$ coincides with $G_{n}$. The conclusion follows. ∎
Let $\hat{\lambda}$ be the law of the UBGW tree of i.i.d weights $\omega$ and reproduction law $\pi$, let $\hat{\lambda}_{n}$ be the measure arising from $G_{n,\infty}^{(i,j)}$ conditionally on $(i,j)\in E_{n}$ and $i,j$ picked uniformly. We also denote by $\hat{\lambda}_{H}$ and $\hat{\lambda}_{n,H}$ the corresponding measures when restricted to the space of trees of height $H$. Lemma 18 states that:
@@
From $(\mathbb{T},\mathbb{M}_{h})$, we define a function on the space of rooted graphs $\hat{\mathcal{G}}^{*}$ in the following way:
$g(T)=\mathbb{P}(o\in\mathbb{M}|\mathbb{T}=T),\forall T\in\hat{\mathcal{G}}^{*},$ (21)
@@
Integrating with regards to the second marginal, we can rewrite $\operatorname{perf_{E}}(\mathbb{T},\mathbb{M}_{h})$ as
$\operatorname{perf_{E}}(\mathbb{T},\mathbb{M}_{h})=\mathbb{E}\left[w(o)\mathbbm{1}_{o\in\mathbb{M}_{h}}\right]=\int_{\hat{\mathcal{G}}^{*}}w(o)g(T)\hat{\lambda}(\mathrm{d}T),$ (22)
@@
$g_{H}(T)=\mathbb{P}(o\in\mathbb{M}_{h}|N_{H}(\mathbb{T})=N_{H}(T)),\forall T\in\hat{\mathcal{G}}^{*}.$ (23)
The function $g_{H}$ can be seen as a conditional expectation if we only looked at $T$ up to depth $H$. The quantity $g_{H}(T)$ is exactly what we wanted at the beginning: it returns a matching score for the root if $\mathbb{M}_{h}$sees" $T$ up to depth $H$. For $x\in\mathbb{R}$, define $g_{H,x}=g_{H}\mathbbm{1}_{w(o)\leq x}$.
@@
We can finally define our random "score" matrix:
@@
$Q_{H,n,x}(G_{n}):=\left(q_{i,j,H,n,x}\right)_{(i,j)\in V_{n}^{2}}$ is a $|V_{n}|\times|V_{n}|$ random matrix with:
$\forall(i,j)\in V_{n}^{2},i\neq j,\,q_{i,j,H,n,x}:=g_{H,x}(G_{n,\infty}^{(i,j)})$
and $q_{j,j,H,n,x}:=1-\sum_{i\neq j}q_{j,i,H,n,x}$ for $j\in V_{n}$.
We will use several intermediary results which proofs are postponed to Subsection 4.2.3. For all $n\in\mathbb{N}$ Let $w_{i,j}$ be the weights of $(i,j)$ in $(V_{n})^{2}$ (we set it as 00 if the edge is not present).
Lemma 19.
@@
As a consequence, since we defined the matrices $Q_{H,n,x}$ so that the rows sum to $1$, the matrices $Q_{H,n,x}$ are asymptotically left stochastic. Let $\text{Birk}_{n}$ be the Birkhoff polytope on $M_{n}(\mathbb{R})$, $\text{Sym}_{n}$ the subspace of symmetric matrices, $\text{RStoc}_{n}$ the right stochastic matrices for which every row sums to 1 and $\text{LStoc}_{n}$ the left stochastic matrices for which every column sums to 1.
Now, Lemma 20 controls the distance of the matrix $Q_{H,n,x}$ to $\text{LStoc}_{|V_{n}|}$. Since $Q_{H,n,x}$ is symmetric, it also controls the distance to $\text{RStoc}_{|V_{n}|}$. Now, we would like to control the distance to $\text{Birk}_{|V_{n}|}$.
For this, we need some geometric results: We will write $\|M\|_{1}=\sum_{i,j}|m_{i,j}|$ for $M_{n}(\mathbb{R})$. We will use the following proposition that states that if a matrix $M$ is $o(n)$ close to $\text{LStoc}_{n}$ and $o(n)$ close to $\text{RStoc}_{n}$, then it is $o(n)$ close to $\text{LStoc}_{n}\cap\text{RStoc}_{n}=\text{Birk}_{n}$:
Proposition 21.
Assume $M$ has non-negative coefficients.
@@
$\sum_{i}\left|\sum_{j}m_{i,j}-1\right|+\sum_{j}\left|\sum_{i}m_{i,j}-1\right|<n\varepsilon$
@@
$d_{\|\cdot\|_{1}}(M,\mathrm{Birk}_{n})\leq C\varepsilon n.$
By Proposition 21, the distance of $Q_{H,n,x}$ to $\text{Birk}_{|V_{n}|}$ is thus $o(|V_{n}|)$. We need another geometric lemma that controls the distance to symmetric bistochastic matrices from the distance to bistochastic matrices and the distance to symmetric matrices:
Lemma 22.
Let $M\in M_{n}(\mathbb{R})$:
$d(M,\mathrm{Sym}_{n}\cap\mathrm{Birk}_{n})\leq d(M,\mathrm{Sym}_{n})+d(M,\mathrm{Birk}_{n}).$ (26)
Combining Equation (26) with Lemmas 19 and 20, the expected distance of $Q_{H,n,x}$ to symmetric bistochastic matrices is $o(|V_{n}|)$ as $n,H\rightarrow\infty$.
4.2.2 Proof of the second part of Proposition 16
@@
Applying Lemma 19 , there exists $H_{\eta}$, $x_{\eta}$ and $N_{\eta}$ big enough such that for any $n\geq N_{\eta}$:
@@
Now, applying what we just showed at the end of the previous subsection, by taking $H_{\eta,\varepsilon}$ and $N_{\eta,\varepsilon}$ big enough, such that for any $n\geq N_{\eta,\varepsilon}$, the previous inequality holds along with:
@@
Applying Markov inequality:
$\mathbb{P}\left[\frac{1}{|V_{n}|}d(Q_{H_{\eta,\varepsilon},n,x_{\eta}},\mathrm{Birk}_{|V_{n}|}\cap\mathrm{Stoc}_{|V_{n}|})\geq\frac{\varepsilon}{x_{\eta}}\right]\leq\varepsilon$
@@
Involution matrices are in bijection with the set of matchings on the complete graph $K_{|V_{n}|}$. Therefore $\hat{Q}_{H,n,x}$ encodes a random matching $M(\hat{Q}_{H,n,x})$ on $G_{n}$ (we just leave edges unmatched if they are not in $G_{n}$). On the event where $\hat{Q}_{H,n,x}$ exists we have:
@@
Taking the limit inferior we get:
@@
In conclusion, we showed that:

### v2

@@ 3.4 Uniqueness
Lemma 3.3 implies that $\mathbb{P}_{Z\sim\zeta_{h}}(Z=0)>0$. This uniform lower bound does not depend on the values of $w(v_{0},v_{-1}),Z_{h}(v_{0},v_{-1})$, so it is $\mathcal{P}$ almost-sure. ∎
@@
$u_{1}(u)=\operatorname*{argmax}_{y\sim u}(w^{s}_{h}(u,y)-Z_{h}^{s}(u,y)),u_{2}(u)=\operatorname*{\overset{[2]}{argmax}}_{y\sim u_{1}}(w^{s}_{h}(u_{1},y)-Z_{h}^{s}(u_{1},y)),$
@@
$u_{2n+1}(u)=\operatorname*{argmax}_{y\sim u_{2n}}(w^{s}_{h}(u_{2n},y)-Z_{h}^{s}(u_{2n},y)),u_{2n+2}(u)=\operatorname*{\overset{[2]}{argmax}}_{y\sim u_{2n+1}}(w^{s}_{h}(u_{2n+1},y)-Z_{h}^{s}(u_{2n+1},y)).$
@@
4 From infinite to finite : Theorem 2
In this entire section, unless stated otherwise, we will adopt the edge-rooted point of view. The aim of this section is to prove Theorem 2, namely that the optimal matching on a random graph converges locally to the unimodular optimal matching on the limiting UBGW tree. Theorem 1, proved in Section 3, states that there is a unique unimodular optimal matching on a UBGW tree with iid weights. A natural approach to prove Theorem 2 is then to show that every subsequential limit of optimal matchings is a unimodular optimal matching on the limiting tree. More precisely, we will prove the following statement:
Proposition 4.1.
Assume $(G_{n},o_{n})_{n\geq 1}$ is a sequence of finite graphs which converges locally to the unimodular BGW tree $(\mathbb{T},o)$ with reproduction law $\pi$ and weights $\omega$.
@@ 4 From infinite to finite : Theorem 2
The proof can be broken down into two parts, one relatively easy, and one more challenging:
@@
The complexity lies in showing that the limit of any convergent subsequence is necessarily optimal. We will do this by constructing an almost optimal matching on $G_{n}$ from the optimal matching $(\mathbb{T},\mathbb{M}_{h})$. Indeed, as a consequence, this will show that $M_{\mathrm{opt}}(G_{n})$ does asymptotically better than $\operatorname{perf_{V}}(\mathbb{M}_{h})$.
@@ 4.1 Local convergence of subsequences
Lemma 4.2.
Let $(G_{n},o_{n})$ be a sequence in the space of rooted graphs, $G_{n}=(V_{n},E_{n})$. Let $(f_{n})_{n\in\mathbb{N}}$ be any (random) functions on $E_{n}$ such that the sequence $(f_{n})$ is uniformly bounded. Assume that $G_{n}$ converges locally to an almost surely locally finite graph $G$. Then for any $H\in\mathbb{N}$, the sequence of decorated graphs $B_{H}(G_{n},o_{n},f_{n})$ is tight.
@@
Let $H\in\mathbb{N}$ and $\varepsilon>0$. Denote by $\tau_{H,N}$ the set of rooted graphs of depth $H$ and degree bounded by $N$. Since $G$ is locally finite, for $N$ large enough, we have $\mathbb{P}(B_{H}(G,o)\in\tau_{H,N})>1-\frac{\epsilon}{2}$. By local convergence, $\mathbb{P}(B_{H}(G_{n},o_{n})\in\tau_{H,N})\underset{n\rightarrow\infty}{\longrightarrow}\mathbb{P}(B_{H}(G,o)\in\tau_{H,N})$ hence for $n$ large enough,
$\mathbb{P}(B_{H}(G_{n},o_{n})\in\tau_{H,N})>1-\epsilon$
@@
$\mathbb{P}(B_{H}(G_{n},o_{n},f_{n})\in\tau_{H,N,A})>1-\epsilon,$
$B_{H}(G_{n},o,f_{n})$ is thus tight. ∎
We can now go back to the proof of Proposition 4.1. Take $f_{n}=\mathbbm{1}_{M_{\mathrm{opt}}(G_{n})}$, it is uniformly bounded by $1$. Apply Lemma 4.2 to $B_{0}(G_{n},o_{n},f_{n})$ to obtain a subsequence such that $B_{0}(G_{k_{n}},o_{k_{n}},f_{k_{n}})$ converges.
But $(G_{k_{n}},o_{k_{n}})$ still converges locally to $(G,o)$, so apply Lemma 4.2 again to find a subsequence $k_{n}^{\prime}$ of $k_{n}$ such that $B_{1}(G_{k^{\prime}_{n}},o_{k^{\prime}_{n}},f_{k^{\prime}_{n}})$ converges locally. Repeat the process and use a diagonal argument to obtain a subsequence $l_{n}$ such that for all $H\in\mathbb{N}$, $B_{H}(G_{l_{n}},o_{l_{n}},f_{l_{n}})$ converges weakly to some limit $B_{H}(G,o,g_{H})$ for some decoration $g_{H}$.
Furthermore, since the decorations $g_{H}$ are obtained from a diagonal extraction, the family $(g_{H})_{H\geq 0}$ is consistent and therefore defines a limiting decoration $g$. To prove that this decoration is a matching, we just have to take successively limits in $n\rightarrow\infty$ and $H\rightarrow\infty$ in the equalities
$\mathbb{P}\left(\sum_{u\sim v}f_{l_{n}}(u,v)\leq 1,(u,v)\in V(B_{H}(G_{n},o_{n}))\right)=1,\mathbb{P}\left(f_{l_{n}}(e)\in\{0,1\},e\in E(B_{H}(G_{n},o_{n}))\right)=1.$
@@ 4.2 Optimality of limits of subsequences
In this section, we carry out the hard part of the proof of Proposition 4.1. We want to reconstruct a quasi matching on $G_{n}=(V_{n},E_{n})$ from $(\mathbb{T},\mathbb{M}_{h})$ in the sense that the quasi matching is a subgraph that differs from a true matching with $o(|E_{n}|)$ amount of edges . The crucial idea is, for every $(i,j)\in V_{n}^{2}$, to compute the probability that $(i,j)\in\mathbb{M}_{h}$ conditionally on the $H$-neighbourhood of $o$ in $\mathbb{T}$ being the corresponding one of $(i,j)$ in $G_{n}$. Informally, this constructs a "score" matrix where each score is the likelihood of matching $(i,j)$ by looking up to depth $H$.
The first hurdle is that the $H$- neighbourhood of $(i,j)$ in $G_{n}$ may not be a tree, so we will consider the universal cover of $G_{n}$ rooted at $(i,j)$ instead (see below for a definition), and control the resulting error.
@@
By construction, this random matching on $G_{n}$ will have the same performance as $(\mathbb{T},\mathbb{M}_{h})$ with asymptotically small error, which yields the result.
@@ 4.2.1 Score function on edges of finite graphs
Let $(i,j)\in V_{n}^{2}$, we define $G_{n,\infty}^{(i,j)}$ as $\emptyset$ if $\{i,j\}$ is not an edge of $G_{n}$, and as the universal cover $G_{n,\infty}^{(i,j)}$ of $G_{n}$ rooted at the oriented edge $(i,j)$ if $\{i,j\}$ is an edge of $G_{n}$. This cover is the tree of non-backtracking trajectories from the edge $(i,j)$. Namely, when $\{i,j\}$ is an edge of $G_{n}$, $G_{n,\infty}^{(i,j)}$ is a tree rooted at the oriented edge $(i,j)$, children of $i$ (resp. $j$) are the neighbours of $i$ (resp. $j$) in $G_{n}$, with $j$ (resp. $i$) excluded. The children of any vertex $v$ with parent $v^{\prime}$ are then the neighbours of $v$ in $G_{n}$, with $v^{\prime}$ excluded. See Figure 14 for an illustration. For $H\geq 0$, we set $G_{n,H}^{(i,j)}:=B_{H}(G_{n,\infty}^{(i,j)})$.
@@
We will need the following statement that translates the fact that the universal cover of $G_{n}$ converges locally to the same limit as $G_{n}$.
Lemma 4.3.
Let $(i,j)$ be chosen uniformly in $E_{n}$, then:
$G_{n,\infty}^{(i,j)}\underset{n\rightarrow\infty}{\overset{loc}{\longrightarrow}}\mathbb{T}.$
Proof of Lemma 4.3.
@@
$\mathbb{P}\left(B_{H}(G_{n},(i,j))\text{ is a tree}\right)\underset{n\rightarrow\infty}{\longrightarrow}\mathbb{P}\left(B_{H}(\mathbb{T})\text{ is a tree}\right)=1.$
So w.h.p $B_{H}((G_{n},(i,j)))$ is a tree, in which case the universal cover of $G_{n}$ rooted on $o_{n}$ up to height $H$ coincides with $G_{n}$. The conclusion follows. ∎
Let $\hat{\lambda}$ be the law of the UBGW tree of i.i.d weights $\omega$ and reproduction law $\pi$, let $\hat{\lambda}_{n}$ be the measure arising from $G_{n,\infty}^{(i,j)}$ with $(i,j)$ picked uniformly in $E_{n}$. We also denote by $\hat{\lambda}_{H}$ and $\hat{\lambda}_{n,H}$ the corresponding measures when restricted to the space of trees of height $H$. Lemma 4.3 states that:
@@
From $(\mathbb{T},\mathbb{M}_{h})$, we define a function on the space of rooted graphs $\hat{\mathcal{G}}^{\star}$ in the following way:
$g(T)=\mathbb{P}(o\in\mathbb{M}|\mathbb{T}=T),\forall T\in\hat{\mathcal{G}}^{\star},$ (24)
@@
Integrating with respect to the second marginal, we can rewrite $\operatorname{perf_{E}}(\mathbb{T},\mathbb{M}_{h})$ as
$\operatorname{perf_{E}}(\mathbb{T},\mathbb{M}_{h})=\mathbb{E}\left[w(o)\mathbbm{1}_{o\in\mathbb{M}_{h}}\right]=\int_{\hat{\mathcal{G}}^{\star}}w(o)g(T)\hat{\lambda}(\mathrm{d}T),$ (25)
@@
$g_{H}(T)=\mathbb{P}(o\in\mathbb{M}_{h}|B_{H}(\mathbb{T})=B_{H}(T)),\forall T\in\hat{\mathcal{G}}^{\star}.$ (26)
The function $g_{H}$, which can be seen as a conditional expectation if we only looked at $T$ up to depth $H$, is exactly what we wanted at the beginning: it returns a matching score for the root if $\mathbb{M}_{h}$sees" $T$ up to depth $H$. Furthermore, it can be computed explicitly from the law $\zeta_{h}$, on $B_{H}(T)$ with an identical construction as in Proposition 3.2 as follows. Given $B_{H}(T)$, draw i.i.d variables of law $\zeta_{h}$ on the $H-$boundary pointing away from the root, then use recursion (3) to define it inside $B_{H}(T)$, then, writing $o=(o_{-},o_{+})$, $g_{H}(T)$ is the probability that $Z(o_{-},o_{+})+Z(o_{+},o_{-})\leq w(o)$.
For integrability reasons, we also exclude large weights, for $x\in\mathbb{R}$, define $g_{H,x}=g_{H}\mathbbm{1}_{w(o)\leq x}$.
@@
We can finally define a random "score" matrix:
@@
For all $(i,j)\in V_{n}^{2},i\neq j$, define
$q_{i,j,H,n,x}:=g_{H,x}(G_{n,\infty}^{(i,j)}),$ (27)
and
$q_{j,j,H,n,x}:=1-\sum_{i\neq j}q_{j,i,H,n,x}.$ (28)
The score matrix is the $|V_{n}|\times|V_{n}|$ random matrix defined by
$Q_{H,n,x}(G_{n}):=\left(q_{i,j,H,n,x}\right)_{(i,j)\in V_{n}^{2}}.$ (29)
Once again, $q_{i,j,H,n,x}$ can be computed from the distribution $\zeta_{h}$, see Figure 15.
We will use several intermediary results which proofs are postponed to Subsection 4.2.3. Let $w_{i,j}$ be the weights of $(i,j)$ in $(V_{n})^{2}$ (we set it as 00 if the edge is not present).
Lemma 4.5.
@@
Let $\text{Birk}_{d}$ be the Birkhoff polytope on $M_{d}(\mathbb{R})$, $\text{Sym}_{d}$ the subspace of symmetric matrices, $\text{RStoc}_{d}$ the right stochastic matrices for which every row sums to 1 and $\text{LStoc}_{d}$ the left stochastic matrices for which every column sums to 1.
Lemma 4.6 controls the distance of the matrix $Q_{H,n,x}$ to $\text{LStoc}_{|V_{n}|}$. Since $Q_{H,n,x}$ is symmetric, it also controls the distance to $\text{RStoc}_{|V_{n}|}$. Now, we would like to control the distance to $\text{Birk}_{|V_{n}|}=\text{LStoc}_{|V_{n}|}\cap\text{RStoc}_{|v_{n}|}$.
For this purpose, we need some geometric result. Write $\|M\|_{1}=\sum_{i,j}|m_{i,j}|$ for $M\in M_{d}(\mathbb{R})$. We will use the following proposition that states that if a matrix $M$ is $o(d)$ close to $\text{LStoc}_{d}$ and $o(d)$ close to $\text{RStoc}_{d}$, then it is $o(d)$ close to $\text{Birk}_{d}$:
Proposition 4.7.
Assume $M\in M_{d}(\mathbb{R})$ has non-negative coefficients.
@@
$\sum_{i}\left|\sum_{j}m_{i,j}-1\right|+\sum_{j}\left|\sum_{i}m_{i,j}-1\right|<d\varepsilon.$
@@
$d_{\|\cdot\|_{1}}(M,\mathrm{Birk}_{d})\leq C\varepsilon d.$
By Proposition 4.7, the distance of $Q_{H,n,x}$ to $\text{Birk}_{|V_{n}|}$ is thus $o(|V_{n}|)$. We need another geometric lemma that controls the distance to symmetric bistochastic matrices from the distance to bistochastic matrices and the distance to symmetric matrices:
Lemma 4.8.
Let $M\in M_{d}(\mathbb{R})$:
$d(M,\mathrm{Sym}_{d}\cap\mathrm{Birk}_{d})\leq d(M,\mathrm{Sym}_{d})+d(M,\mathrm{Birk}_{d}).$ (32)
Combining Equation (32) with Lemmas 4.5 and 4.6, the expected distance of $Q_{H,n,x}$ to symmetric bistochastic matrices is $o(|V_{n}|)$ as $n,H\rightarrow\infty$.
4.2.2 Proof of the second part of Proposition 4.1
@@
Applying Lemma 4.5 , there exists $H_{\eta}$, $x_{\eta}$ and $N_{\eta}$ big enough such that for any $n\geq N_{\eta}$,
@@
Now, applying what we just showed at the end of the previous subsection, by taking $H_{\eta,\varepsilon}$ and $N_{\eta,\varepsilon}$ big enough, such that for any $n\geq N_{\eta,\varepsilon}$, the previous inequality holds along with
@@
Applying Markov inequality,
$\mathbb{P}\left[\frac{1}{|V_{n}|}d(Q_{H_{\eta,\varepsilon},n,x_{\eta}},\mathrm{Birk}_{|V_{n}|}\cap\mathrm{Stoc}_{|V_{n}|})\geq\frac{\varepsilon}{x_{\eta}}\right]\leq\varepsilon.$
@@
Involution matrices are in bijection with the set of matchings on the complete graph $K_{|V_{n}|}$. Therefore $\hat{Q}_{H,n,x}$ encodes a random matching $M(\hat{Q}_{H,n,x})$ on $G_{n}$ (we just leave edges unmatched if they are not in $G_{n}$). On the event where $\hat{Q}_{H,n,x}$ exists, we have
@@
Taking the limit inferior we get
@@
In conclusion, we showed that